OpenStudy (anonymous):

evaluate the limit and justify each step by indicating the appropriate limit laws as x approaches 2. lim ((2x^2+1)+1)/(x^2+6x-4)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}=((2x ^{2}+1)+1)\div( x ^{^{2}}+6x-4)\] \[\lim_{x \rightarrow 2}=((2(2) ^{2}+1)+1)\div( 2 ^{^{2}}+6(2)-4)\] \[\lim_{x \rightarrow 2}=10\div12\]=5/6

OpenStudy (anonymous):

thank you so much! i thought that was what it was but i wasn't sure!